SHOW ALL WORK!!!!!

Ameth Growers has historically had a P/E ratio of 21.4. This ratio is considered a good estimate of the future ratio. The firm currently has EPS of $2.34. These earnings are expected to increase by 3.4 percent next year. What is the expected price of this stock one year from now?

SHOW ALL WORK!!!!!
Answer #1

EPS next year = current EPS*(1+growth rate) = 2.34*(1+0.034)=2.41956

Price = P/E*EPS next year = 21.4*2.41956=51.78